OpenStudy (mathmath333):

An amount of Rs. 12820 due 3 years hence, is fully repaid in three annual installments staring after 1 year. The first second and third installments are in ratio 1:2:3. If the rate of compound interest is 10% per annum, find the first installment

OpenStudy (mathmath333):

a. 2400 b. 1800 c. 2000 d. 2500

OpenStudy (mathmath333):

it has straight forward solution of \(\large\tt \begin{align} \color{black}{ a(1.1^²) + 2a(1.1) + 3a = 12820 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

still looks strange to me

OpenStudy (mathmath333):

@freckles

OpenStudy (mathmath333):

@ganeshie8

ganeshie8 (ganeshie8):

*

OpenStudy (mathmath333):

its \(\huge a\)

OpenStudy (freckles):

and you squared that one 1.1 because we were in the third year

OpenStudy (freckles):

I could be completing wrong but your equation seems fine to me

OpenStudy (freckles):

completely*

ganeshie8 (ganeshie8):

still feeling stuck on this @mathmath333 ?

OpenStudy (mathmath333):

yes still stuck @ganeshie8

OpenStudy (mathmath333):

a sum of 8000 is borrowed at 5% p.a. compound interest and paid back in 3 equal annual instalments .what is the amount of each installment?\(\large\tt \begin{align} \color{black}{\dfrac{x}{(1+\dfrac{5}{100})^1}+\dfrac{x}{(1+\dfrac{5}{100})^2}+\dfrac{x}{(1+\dfrac{5}{100})^3}=8000 \hspace{.33em}\\~\\ }\end{align}\) for the \(\Huge \uparrow\) question the rate % is in denominator while for the posted question it is in nummerator

ganeshie8 (ganeshie8):

did you watch the movie back to the future

ganeshie8 (ganeshie8):

In the first question the debt is due in future In the second question, you took the loan already so the debt is due from today!

ganeshie8 (ganeshie8):

First question is exactly same as below : find the FUTURE VALUE of money if you make payments each year. first year an amount of \(a\), second year an amount of \(2a\) and in the third year an amount of \(3a\)

ganeshie8 (ganeshie8):

assuming payments are made at the end of each year, notice that the first payment stays in bank for exactly 2 years and the interest compounds 2 times : so the FUTURE VALUE of first payment of \(a\) would be : \(a(1+\frac{10}{100})^2 = a(1.1)^2\)

OpenStudy (mathmath333):

in the first question \(12820\) is compounded amount or the principle

ganeshie8 (ganeshie8):

12820 is the compounded FUTURE amount ! there is no starting principal in the first question

OpenStudy (mathmath333):

ohk i get that future value formula

OpenStudy (mathmath333):

oh lol i was assuming it as the principle haha

OpenStudy (mathmath333):

it was clearly stated as " \(\bf amount\) "of Rs \(12820\) still i was beaten

ganeshie8 (ganeshie8):

Ahh I see... that first question is equivalent to the problem of finding the FUTURE VALUE of a savings account where the payments are made at regular intervals. each payment earns a different interest because of the time value of money

ganeshie8 (ganeshie8):

a $100 TODAY is worth more than $1000 after 10 years

OpenStudy (mathmath333):

so in the previously derived formula it should \(\large\tt \begin{align} \color{black}{x+\left(x+\dfrac{x\times r\%\times 1}{100}\right)+\left(x+\dfrac{x\times r\%\times 2}{100}\right)\\\cdot \cdot \cdot \cdot \cdot \left(x+\dfrac{x\times r\%\times (n-1)}{100}\right) =\cancel P~~\color{red }{A}\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

it should \(A\) not \(P\)

ganeshie8 (ganeshie8):

thats right!

OpenStudy (mathmath333):

also for this one \(\large\tt \begin{align} \color{black}{\dfrac{x}{(1+\dfrac{r}{100})^1}+\dfrac{x}{(1+\dfrac{r}{100})^2}\\+\cdot \cdot \cdot \cdot \cdot \cdot \dfrac{x}{(1+\dfrac{r}{100})^n}=\cancel P \color{blue}{A} \hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

Looks good! in the first question you will be `earning` interest because it is same as saving money in your bank. in the second question you will be `paying` the interest to bank because you took the loan today.

OpenStudy (mathmath333):

ohk that helps now

OpenStudy (mathmath333):

thanks!